OpenStudy (anonymous):

ln(x)+ln(x-1)=1 so far ive completed the square and now have: (x-1/2)^2=e+1/4

OpenStudy (lgbasallote):

what happened?

OpenStudy (anonymous):

what i think im supposed to do isnt what the book is showing i would square both sides next and get x-1/2=sqrt(e+1/4) then add the 1/2 x=1/2+sqrt(e+1/4)

OpenStudy (lgbasallote):

start by applying product law

OpenStudy (lgbasallote):

apply product rule on ln x + ln (x-1) what do you get?

OpenStudy (anonymous):

ln(x(x-1))

OpenStudy (lgbasallote):

good so.. \[\ln [x(x+1)] = 1\] now... i can change 1 into ln (0) agree?

OpenStudy (lgbasallote):

uhh nevermind

OpenStudy (lgbasallote):

change to exponential form

OpenStudy (anonymous):

so e^1

OpenStudy (anonymous):

he's already done all that.

OpenStudy (lgbasallote):

right x(x+1) = e

OpenStudy (lgbasallote):

really?

OpenStudy (anonymous):

yes but continue

OpenStudy (lgbasallote):

my bad

OpenStudy (anonymous):

x(x-1) = e anyway

OpenStudy (anonymous):

i dont mind lets do this step by step

OpenStudy (lgbasallote):

distribute x into x-1

OpenStudy (anonymous):

x^2-x=e

OpenStudy (anonymous):

you're done... take the root of both sides and simplify...

OpenStudy (lgbasallote):

now use completing the square method to get the third term in x^2 - x

OpenStudy (anonymous):

(x-1/2)^2=e+1/4

OpenStudy (lgbasallote):

now take the square root of both sides

OpenStudy (anonymous):

x-1/2=sqrt(e+1/4)

OpenStudy (lgbasallote):

wrong,.. \[x - \frac 12 = \pm \sqrt{e+\frac 14}\] that plus-minus is important

OpenStudy (anonymous):

lol correct my bad

OpenStudy (lgbasallote):

so now add 1/2

OpenStudy (anonymous):

andddd i think i see where i went wrong lol

OpenStudy (lgbasallote):

the plus-minus?

OpenStudy (anonymous):

x=1/2+_ sqrt(e+1/4)

OpenStudy (lgbasallote):

that's not yet simplified though

OpenStudy (lgbasallote):

express e + 1/4 as one fraction

OpenStudy (anonymous):

(4e+1)/4?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now... simplify \[\sqrt{\frac{4e+1}4}\]

OpenStudy (anonymous):

\[x=\frac{ 1 }{ 2 }\pm \sqrt{\frac{ 4e+1 }{ 4 }}\]

OpenStudy (lgbasallote):

do you now the square root of 4?

OpenStudy (anonymous):

x=1/2+-(sqrt (4e+1))/2

OpenStudy (lgbasallote):

right. now express as one fraction

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(1\pm \sqrt{4e+1})\]

OpenStudy (lgbasallote):

right. and that is how mathematicians solve problems. with meticulousnessand OCD

OpenStudy (anonymous):

lol ive been on that problem for like 20 min thanks

OpenStudy (lgbasallote):

just doing my part in helping an awesome user in beating satellite73's SmartScore